bivibes:

Andrea has a jug containing 1.5 liters of chocolate milk. She fills one cup with 250 milliliters and another cup with 0.6 liters. How much chocolate milk is left in the jug in milliliters? 0.65 milliliters 249.1 milliliters 450 milliliters 650 milliliters

bivibes:

450?

AZ:

You need to subtract 1.5 liters - 0.6 liters - 250 milliliters First, make them all have the same units How many milliliters is 1.5 liters? and 0.6 liters is how many milliliters?

AZ:

@bivibes wrote:
450?
How did you get that?

bivibes:

bc

bivibes:

idk ;-;

AZ:

How many milliliters is 1.5 liters? We just answered this in your previous question

bivibes:

1500

AZ:

Good and how many milliliters are there in 0.6 Liters? Remember, 1 liter has 1000 milliliter so 0.6 liters would have 0.6 * 1000 = ??? milliliters

bivibes:

2100?

AZ:

What is 0.6 * 1000 = ?

AZ:

It is not 2100 Use a calculator if you must

bivibes:

600

bivibes:

-_-

bivibes:

for sum dum reason i added it with 1500 :?

AZ:

Yes Now look back at the original question You had a jug with 1500 milliliters of chocolate milk You fill one cup with 250 milliliters You fill another cup with 600 milliliters How much milk is left? 1500 - 250 - 600 = ??

bivibes:

650

AZ:

And there you go! That's your answer in terms of milliliters

bivibes:

oml tysm

AZ:

No problem